# references.bib

### v2

@@ 1.1 Contributions
  • how torsion is typically lost due to dimensionality reduction in the encoder (Section 5.2),
  • how torsion that is preserved in the latent space leads to changes in persistence diagrams over field coefficients (Section 6),
  • and which decoder architectures fail to reconstruct torsion — and in which cases reconstruction becomes possible (Section 7.1).
@@ 2 Persistent Homology
[…]
@@
on the set of simplices called a filtration. In the following, we work with Vietoris-Rips filtrations (in short: “Rips filtrations”).
Definition 2.1 (Vietoris-Rips Filtration).
[…]
@@
We assume in the following that filtrations are embedded in $\mathbb{R}^{d}$, thereby excluding phenomena like Vietoris-Rips bubbles [vrplanarpoint]. The $p$-th persistent homology of a finite-dimensional filtered simplicial complex with coefficients in a commutative ring $R$ is given by
@@
where $H_{p}(K_{i},R)$ denotes the simplicial homology group for a subcomplex $K_{i}$, $0\leq i\leq N$ in the filtration. This sequence captures how topological features such as connected components, loops, and voids appear and disappear across different scales. The image of the map $H_{p}(K_{i};R)\to H_{p}(K_{j};R)$ is referred to as a persistent homology group and identifies features that persist across the filtration interval $[i,j)$.
@@ 2.1 Field Coefficients and Persistence Diagrams
For field coefficients, the torsion subgroup vanishes, and persistent homology is fully determined by the ranks of the free parts of the persistent homology groups (field Betti numbers). As a result, its topological invariants admit a simple representation: namely, the persistence intervals $[i,j)$ over which a feature exists within the filtration encode all relevant information. These intervals can be visualized either as points in a persistence diagram—where the $x$-axis denotes the birth time of a feature and the $y$-axis its death time—or as horizontal bars in a so-called barcode diagram. In the following, we denote by $D_{p}(X;R)$ the persistence diagram of a topological space $X$ in homological dimension $p$, with coefficients in a ring $R$.
@@ 2.2 Torsion
Torsion coefficients annihilate specific elements of the homology group via multiplication. Specifically, for each element of the torsion subgroup $z\in H_{p}(K;\mathbb{Z})$, there exists an integer $q^{k}$, for $q$ prime, such that
@@
Let $X$ be a topological space, $k$ a coefficient field and $p$ a homology dimension. The persistence diagram $D_{p}(X;k)$ is independent of the choice of coefficient field, if the relative integral homology group $H_{p}(X_{i},X_{j};\mathbb{Z})$ is free for any indices $0\leq i<j\leq N$ in the filtration of $X$ and $H_{p-1}(X_{i};\mathbb{Z})$ is free for any $0\leq i\leq N$.
@@ 3 Neural Networks
[…]
@@ 4 Autoencoders
With regard to their broad applicability and the need for non-model driven ways for data analysis, there exists a wide variation of autoencoder architectures. We call an autoencoder trained using the MSE loss vanilla. In Appendix B, we present two modifications of the loss term inspired from TDA, that we use for subsequent experiments in this work.

### v3

# Torsion in Persistent Homology and Neural Networks

@@ 1.1 Contributions
  • how torsion is typically lost due to dimensionality reduction in the encoder (Section LABEL:sec:dimred),
  • how torsion that is preserved in the latent space leads to changes in persistence diagrams over field coefficients (Section LABEL:sec:torsreplatsp),
  • and which decoder architectures fail to reconstruct torsion — and in which cases reconstruction becomes possible (Section LABEL:sec:lindec).
@@ 2 Persistent Homology
We define persistent homology with regard to its application to machine learning settings. In this context, we understand a data set $\mathcal{D}=\{x_{i}\in\mathbb{R}^{d}\;|\;1\leq i\leq n,\;n\in\mathbb{N}\}$ as a finite (Euclidean) metric space, also called a point cloud. On this data set we construct a finite geometric simplicial complex – a topological space constructed by gluing together dimensional simplices (points, line segments, triangles, tetrahedrons, etc.). We impose a total order
@@
on the set of simplices called a filtration. In the following, we work predominantly with Vietoris–Rips filtrations (in short: “Rips filtrations”).
Definition 2.1 (Vietoris–Rips Filtration).
Given a point cloud $\mathcal{D}=\{x_{i}\in\mathbb{R}^{d}\;|\;1\leq i\leq n\}$} and positive real numbers $\delta_{0}<\delta_{1}<\cdots<\delta_{m}$ with $\delta_{i}\in\mathbb{R}_{>0}$, the Vietoris–Rips filtration is the nested sequence of simplicial complexes
@@
Definition 2.1 introduces the notion of an abstract filtered simplicial complex, a purely combinatorial object that can be endowed with the structure of a topological space via geometric realization ([edelharerbook], Section III.1). For a Vietoris–Rips filtration constructed on a finite metric space $M\subset\mathbb{R}^{d}$, the dimension of its geometric realization may exceed $d$. In the following, we restrict attention to geometric filtered simplicial complexes that are embedded in $\mathbb{R}^{d}$, thereby excluding phenomena such as Vietoris–Rips bubbles [vrplanarpoint]. Our objects of study are thus finite geometric filtered simplicial complexes $K$ constructed from finite metric spaces $M\subset\mathbb{R}^{d}$. The $p$-th persistent homology of a finite-dimensional filtered simplicial complex with coefficients in a commutative ring $R$ is given by
@@
where $H_{p}(K_{i},R)$ denotes the $p$-th simplicial homology group for a subcomplex $K_{i}$, $0\leq i\leq N$, in the filtration. This sequence captures how topological features such as connected components, loops, and voids appear and disappear across different scales. The image of the map $H_{p}(K_{i};R)\to H_{p}(K_{j};R)$ is referred to as a persistent homology group and identifies features that persist across the filtration interval $[i,j)$.
@@ 2.1 Field Coefficients and Persistence Diagrams
For field coefficients, the torsion subgroup vanishes, and persistent homology is fully determined by the ranks of the free parts of the persistent homology groups (field Betti numbers). As a result, its topological invariants admit a simple representation: namely, the persistence intervals $[i,j)$ over which a feature exists within the filtration encode all relevant information. These intervals can be visualized either as points in a persistence diagram—where the $x$-axis denotes the birth time of a feature and the $y$-axis its death time—or as horizontal bars in a so-called barcode diagram. In the following, we denote by $\mathrm{dgm}_{p}(X;R)$ the persistence diagram in homological dimension $p$, with coefficients in the ring $R$, of a topological space $X$ obtained as the topological realization of a filtration of a finite metric space.
@@ 2.2 Torsion
Torsion coefficients annihilate specific elements of the homology group via multiplication. Specifically, for each element $z\in H_{p}(K;\mathbb{Z})$ of the torsion subgroup, there exists an integer $q^{k}$, for $q$ prime, such that
@@
Let $X$ be a topological space obtained as the topological realization of a filtration of a finite metric space, $k$ a coefficient field and $p$ a homology dimension. The persistence diagram $\mathrm{dgm}_{p}(X;k)$ is independent of the choice of coefficient field, if the relative integral homology group $H_{p}(X_{i},X_{j};\mathbb{Z})$ is free for any indices $0\leq i<j\leq N$ in the filtration of $X$ and $H_{p-1}(X_{i};\mathbb{Z})$ is free for any $0\leq i\leq N$.
@@ 3 Neural Networks
Let $\mathcal{D}=\{x_{i}\in\mathbb{R}^{d}\;|\;1\leq i\leq n,\;n\in\mathbb{N}\}$ be a data set serving as the input to the neural network. The output $o_{ijk}$ of neuron $k$ in layer $j$ for an input $x_{i}\in\mathbb{R}^{v}$ is computed as:
@@ 4 Autoencoders
With regard to their broad applicability and the need for non-model driven ways for data analysis, there exists a wide variation of autoencoder architectures. We call an autoencoder trained using the MSE loss vanilla. In Appendix LABEL:app:AE, we present two modifications of the loss term inspired from TDA, that we use for subsequent experiments in this work.